OpenStudy (anonymous):

What are the exact solutions of x2 - 5x - 7 = 0?

OpenStudy (anonymous):

Do you understand ot are you completely lost?

OpenStudy (anonymous):

iam completely lost

OpenStudy (anonymous):

well do you know what exact solutions are?

OpenStudy (anonymous):

solutions to the equation I think

OpenStudy (anonymous):

well sorta (want me to explain?)

OpenStudy (anonymous):

yeah please

OpenStudy (anonymous):

As used in physics, the term "exact" generally refers to a solution that captures the entire physics and mathematics of a problem as opposed to one that is approximate, perturbative, etc. Exact solutions therefore need not be closed-form.

OpenStudy (anonymous):

should I put up the answers

OpenStudy (anonymous):

yes plz do

OpenStudy (anonymous):

hey hold on ill be right back im bout to change my picture i look old in this one

OpenStudy (anonymous):

x = the quantity of negative 5 plus or minus the square root of 3 all over 2 x= the quantity of 5 plus or minus the square root of 3 all over 2 x = the quantity of negative 5 plus or minus the square root of 53 all over 2 x = the quantity of 5 plus or minus the square root of 53 all over 2

OpenStudy (anonymous):

x=\[x={\frac{ 5 \pm \sqrt(53) }{ 2 }}\]

OpenStudy (anonymous):

howd u get that

OpenStudy (anonymous):

if equation is \[ax^2+bx+c=0\] then solutions of it iare \[\frac{ -b \pm \sqrt(b^2-4ac) }{ 2a }\]

OpenStudy (anonymous):

oh ok I kinda get it now

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thank you